11.      Consider in perfectly competitive market the following demand and supply equations for sugar:

Qd =1000-1000p        where Q d is quantity demanded and Qs is quantity supplied. Qs=800+ 1000p

Where P is the price of sugar per pound and Q is thousands of pounds of sugar.

a.  What are the equilibrium price and quantity for sugar? (1point)

b.  Suppose that the government wishes to subsidize sugar production by placing a floor on sugar prices of $0.20 per pound. What would be the relationship between the quantity supplied and quantity demand for sugar?

C. Identify market problem specifically at prices 0.2 per pound and what will be scientific recommendation you suggest to solve the identified market problem?

 

1
Expert's answer
2021-04-22T09:29:13-0400

a)


"1000-1000p=800+1000p"

"p=0.1"

"Q=900"

b)


"Q_s=800+1000(p+0.2)"

c)


"1000-1000p=800+1000p+200"

"p=0"

There is no need to provide subsidies.